A 55.0m wire carries a current of 2.25 A in a location where Earth's magnetic field is 4.5 x10-3 T. Ther wire experiences a force of 0.15 N. Find the angle between the wire and the magnetic field.


1
Expert's answer
2022-03-28T11:02:04-0400

The Lorentz force

"F=BIl\sin\theta"

Hence, the angle between the wire and the magnetic field

"\theta=\tan^{-1}\frac{F}{BIl}"

"\theta=\tan^{-1}\frac{0.15}{4.5*10^{-3}*2.25*0.55}=88^\circ"